# Leveraging the true depth of LLMs

### v1

@@ 7 Conclusion
In this work, we presented Layer Parallelism, a novel approach that exploits independence patterns between transformer layers to optimize LLM inference. By restructuring the computational graph to enable parallel execution of consecutive layer pairs through tensor parallelism, we achieved substantial speed improvements without model retraining. Our method reduced the effective depth of Llama 2 7B by 21% while maintaining strong performance, yielding up to a 1.29x improvement in inference speed for single-token generation with long sequences. Similar benefits were observed with Llama 3.2 3B, achieving an 18% reduction in effective depth with up to a 1.22x speed-up. Moreover, we show that we can recover 10.8% of ICL accuracy on MMLU by fine-tuning the parallelized models using few resources.
@@
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Ablation: Memory efficiency
Figure 8: Maxiumum memory usage (Mb) to complete the following inference tasks: KV Cache pre-filling for a given sequence length, autoregressive generation up to the indicated sequence length, and single token generation with a pre-filled KV Cache of the indicated sequence length. The baseline is the original model with all layers making use of Tensor Parallelism. The Parallel Layers number ($\Delta$) indicates how many layers have been merged using Layer Parallelism (e.g. a $\Delta$ of 4 indicates that 2 groups of 2 layers have been converted to 2 effective layers).
Appendix B Ablation: Tokens per second
@@
Appendix C Generalization to multiple GPUs
@@
Figure 10: Generalization of Layer Parallelism and Tensor Parallelism. In this case, the figure shows the case for 4 GPUs and 2 layers. The stacked layers represent the tensor parallelism, and the colors indicate the processing of different previously contiguous layers. LP builds on top of TP by assigning the heads from consecutive layers into different GPUs. If there are more GPUs than layers, then each layer will be accelerated using TP, and $Q,K,V\text{ and }att\in\mathbb{R}^{T\times\frac{2D}{g}}$, where $D$ is the feature dimension and $g$ is the total number of GPUs.

### v2

@@ 7 Conclusion
In this work, we presented Layer Parallelism, a novel approach that exploits independence patterns between transformer layers to optimize LLM inference. By restructuring the computational graph to enable parallel execution of consecutive layer pairs through tensor parallelism, we achieved substantial speed improvements without model retraining. We show that we can recover ICL accuracy by fine-tuning the parallelized models.
@@
This paper presents work whose goal is to improve inference serving of LLMs by trading accuracy. Well-intentioned users could find themselves receiving incorrect predictions that could lead to undesirable outcomes.
Acknowledgments
We thank Benjamin Rio for his insightful conversations and valuable input during the preparation of this work. His perspectives influenced several key aspects of the presentation.
Appendix A Ablation: Tokens per second
@@
Appendix B Generalization to multiple GPUs
@@
Figure 9: Layer Parallelism in the case of parallelizing two layers over four accelerators. The stacked layers represent the tensor parallelism, and the colors indicate the processing of different previously contiguous layers. $Q,K,V\text{ and }att\in\mathbb{R}^{T\times\frac{2D}{g}}$, where $D$ is the feature dimension and $g$ is the total number of accelerators.
Layer Parallelism allows one to allocate $N\geq 1$ accelerators for each layer. The implementation remains the same, but now each layer is parallelized using tensor parallelism over its assigned accelerators. Note that both reduction operations (tensor parallel and layer parallel) are nicely executed with a single all-reduce call.
Appendix C Acceleration source
(a) Flame chart of running two standard Tensor-Parallel LLama decoder layers.
(b) Flame chart of running two Llama 3 decoder layers with our Layer Parallelism approach.
Figure 10: Comparison of Flame Graphs when running two consecutive Llama 3.2 3B decoder layers with vanilla tensor parallelism (Fig. 10(a)), and our Layer Parallelism approach (Fig. 10(b)). Note that the time axis scale is different between both graphs. These results were obtained on a workstation using x2 RTX 4090s.
Figure 10 illustrates flame graphs comparing two consecutive Llama 3.2 3B decoder layers using vanilla tensor parallelism and our LP approach. The profiling data summarized in Table 3 reveals that the primary source of acceleration in our LP method stems from reducing the total number of all reduce synchronization operations across GPUs. Specifically, the vanilla tensor parallel approach performs synchronization at every decoder layer, resulting in higher cumulative synchronization overhead. In contrast, our Layer Parallelism implementation runs pairs of layers simultaneously, effectively halving the number of synchronization points. This reduction in synchronization leads to a significant drop in synchronization time from 100.8ms to 50.7ms, directly contributing to the observed improvement in inference speed.
Additionally, Layer Parallelism enables fusion of certain computation kernels—particularly attention and MLP operations across parallelized layers—which further marginally reduces the computation time from 217ms to 208.7ms. Although these computational gains are modest compared to the savings achieved through fewer synchronization operations, kernel fusion further optimizes hardware utilization and enhances overall throughput.
Table 3: Profiling results comparing vanilla Tensor Parallel and Layer Parallel implementations on two consecutive Llama 3.2 3B decoder layers.
Approach Total Time (ms) Sync Time (ms) Computation Time (ms)
Tensor Parallel 317.8 100.8 217.0
Layer Parallel (Ours) 259.4 (x1.23) 50.7 (x1.99) 208.7 (x1.04)